Find the equilibrium of L and K subject to a given output:

minimize cost: C=40K+20L subject to Q=60LK = 3000
1
Expert's answer
2018-02-10T06:23:08-0500
C=40K+20L subject to Q=60LK = 3000
Q=60LK = 3000 , LK =50, then C = 40K + 20L = 40*5 + 20*10=400, L = 10, K=5